OpenStudy (anonymous):

The sum of all angles in a triangle is always 180 degrees. The measure of the first angle is four times the measure of the second angle. The measure of the third angle is 78° more than the second angle. What is each angle?

OpenStudy (anonymous):

Ready to work on this together?

OpenStudy (anonymous):

aw yeah big answer time!

OpenStudy (anonymous):

Awww yeah! Lol. Boog! Still here? We're going to solve this together, but I need your input.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Awesome. First things first, how many degrees will the total equal? As in, if we have: side one + side two + side three = ?????, what is ?????

OpenStudy (anonymous):

i am assuming 180

OpenStudy (anonymous):

That's right! So we have: blahblahblah = 180 So far so good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Awesome. Now let me rewrite the question so we don't have to keep scrolling.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The sum of all angles in a triangle is always 180 degrees. The measure of the first angle is four times the measure of the second angle. The measure of the third angle is 78° more than the second angle. What is each angle?

OpenStudy (anonymous):

Ok. So we see that angle one and angle three are using angle two as our reference point, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

By reference, I mean they're comparing the two angles to the second angle. Does that part make sense?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

Do we know exactly how big angle 2 is right now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Good. So how do we translate this: "The measure of the first angle is four times the measure of the second angle." to math language?

OpenStudy (anonymous):

In other words, angle 1 = "what"?

OpenStudy (anonymous):

thats where i get lost

OpenStudy (anonymous):

Ok, not a problem.

OpenStudy (anonymous):

Think of it this way: If I said "bob is 100 feet tall, and jill is twice as tall", how would you write that as an equation?

OpenStudy (anonymous):

x=100*2

OpenStudy (anonymous):

good! So here's our problem: In our question, we don't know how big angle 2 is. So how do we say "four times bigger than an unknown angle"?

OpenStudy (anonymous):

Or, more directly, how do we say "four times an unknown"?

OpenStudy (anonymous):

x=4*x?

OpenStudy (anonymous):

Well, y = 4x Where y is angle 1, and x is angle 2.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Otherwise, we'd be saying "X is four times bigger than X" which wouldn't really work. ;)

OpenStudy (anonymous):

i see that now

OpenStudy (anonymous):

Alright. So here's what we know so far: Angle one is equal to 4x And angle two is equal to just x.

OpenStudy (anonymous):

The sum of all angles in a triangle is always 180 degrees. The measure of the first angle is four times the measure of the second angle. The measure of the third angle is 78° more than the second angle. What is each angle?

OpenStudy (anonymous):

Show me how we represent angle 3 now.

OpenStudy (anonymous):

And don't let the degree sign throw you off. Just treat it as a 78

OpenStudy (anonymous):

i have no clue

OpenStudy (anonymous):

Well, how do they describe angle three to us? Tell me what they say.

OpenStudy (anonymous):

The measure of the third angle is 78° more than the second angle.

OpenStudy (anonymous):

x=3*78?

OpenStudy (anonymous):

Close. Remember it's saying it's exactly 78 degrees more. Not 78 *times* bigger.

OpenStudy (anonymous):

also, the three is not needed. Even though it's the third angle, we don't use the three for the mathematical expression.

OpenStudy (anonymous):

The measure of the third angle is 78° more than the second angle.

OpenStudy (anonymous):

what was the second ?

OpenStudy (anonymous):

Remember this part: "Alright. So here's what we know so far: Angle one is equal to 4x And angle two is equal to just x."

OpenStudy (anonymous):

x+78?

OpenStudy (anonymous):

i have never had algebra before, so word problems are hard for me to understand

OpenStudy (anonymous):

Perfect! You're absolutely right.

OpenStudy (anonymous):

Oh dear. They should have had you take algebra before you took geometry.

OpenStudy (anonymous):

Or is this algebra class right now?

OpenStudy (anonymous):

collage math 101

OpenStudy (anonymous):

so 4x+78?

OpenStudy (anonymous):

Ah. That is actually algebra 1 then. :D Here we go: you know that: angle 1 is 4x angle 2 is x angle 3 is x+78

OpenStudy (anonymous):

Not yet, I'm afraid. We know from earlier that we said: angle 1 plus angle 2 plus angle 3 = 180 right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so 180 needs to go some where

OpenStudy (anonymous):

if i can figure out the equation i can get the answer

OpenStudy (anonymous):

4x+x+78=180?

OpenStudy (anonymous):

Perfect! Do you know about combing like terms and everything else?

OpenStudy (anonymous):

Oh wait! We forgot ONE thing.

OpenStudy (anonymous):

4x = angle 1 x+78 is angle 3 and x is angle 2 Do you see what you're missing?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hint: putting the expressions in parentheses makes it more obvious: (4x) + (x+78) =180 So what are we missing?

OpenStudy (anonymous):

4x+x+78*2?

OpenStudy (anonymous):

the two didn't come from any of our equations.

OpenStudy (anonymous):

sorry read it wrong

OpenStudy (anonymous):

Hint: (4x) + (x+78) =180 <-- total angles ^ Angle 1 ^ angle 3

OpenStudy (anonymous):

No problem. So you see what's missing, right?

OpenStudy (anonymous):

angle 2

OpenStudy (anonymous):

Perfect!

OpenStudy (anonymous):

now i am lost again

OpenStudy (anonymous):

So we just sneak in an extra +x, right?

OpenStudy (anonymous):

4x + x+78 + x =180

OpenStudy (anonymous):

(4x)+(x+78)+x=180

OpenStudy (anonymous):

Perfect! Now combine like terms.

OpenStudy (anonymous):

(you can now drop the parentheses if you want)

OpenStudy (anonymous):

78 + 4x + x + x = 180

OpenStudy (anonymous):

4x + x = 5x

OpenStudy (anonymous):

78 + 5x + x = 180

OpenStudy (anonymous):

Good! And keep going.

OpenStudy (anonymous):

5x + x = 6x

OpenStudy (anonymous):

78 + 6x = 180

OpenStudy (anonymous):

Excellent! Now what do we do?

OpenStudy (anonymous):

78 + -78 + 6x = 180 + -78

OpenStudy (anonymous):

Absolutely right.

OpenStudy (anonymous):

6x = 180 + -78

OpenStudy (anonymous):

yep. So 6x = 102

OpenStudy (anonymous):

180 + -78 = 102

OpenStudy (anonymous):

And x equals how much?

OpenStudy (anonymous):

17

OpenStudy (anonymous):

so 17 is 1 answer?

OpenStudy (anonymous):

Exactly! Which angle did we decide was 17? As in, what angle did we decide would be "x"?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Perfect! You now have all the information needed to solve.

OpenStudy (anonymous):

The sum of all angles in a triangle is always 180 degrees. The measure of the first angle is four times the measure of the second angle. The measure of the third angle is 78° more than the second angle. What is each angle?

OpenStudy (anonymous):

(4x)+x+78)+17=180?

OpenStudy (anonymous):

well, we only needed that equation to find what x equals. We know that 4*17 + 17+78 + 17 = 180

OpenStudy (anonymous):

Oh crap nevermind that.. i read the question wrong

OpenStudy (anonymous):

Woogy: Tell me mathematically how you got that.

OpenStudy (anonymous):

Oh ok. Mistakes happen. ^_^